OpenStudy (anonymous):

Solve the system by substitution. x - 3y = 4 2x - 6y = 8

OpenStudy (anonymous):

Multiply the first equation by -2: -2x + 6y = -8 2x - 6y = 8 Add the two new equations together. 0 = 0 This is a true statement, so the system has infinitely solutions in the form of x-3y=4. The solution set for this system is a line with that same equation.